OpenStudy (anonymous):

A survey showed that 62% of the people surveyed prefer a certain brand of toothpaste. What fraction of these people prefer the brand of toothpaste?

pooja195 (pooja195):

62% : original = 62/100 : put over 100 as fraction = 31/50 : reduce